OpenStudy (lastdaywork):

Please check my reasoning- For xy>1 \[\tan^{-1}x+\tan^{-1}y=\tan^{-1}\frac{ x+y }{ 1-xy }+\pi\] But by substituting x=(-a) & y=(-b) \[\tan^{-1}a+\tan^{-1}b=\tan^{-1}\frac{ a+b }{ 1-ab }-\pi\] and still ab>1 Hence, the general formula should be \[\tan^{-1}x+\tan^{-1}y=\tan^{-1}\frac{ x+y }{ 1-xy }+sgn(x)* \pi\] for xy>1

OpenStudy (lastdaywork):

PS: I never saw the following formula in books (hence, the need to confirm ;) \[\tan^{-1}x+\tan^{-1}y=\tan^{-1}\frac{ x+y }{ 1-xy }-\pi\] for xy>1

OpenStudy (ikram002p):

so if u dint see it in books did u wrote it by urself ?

OpenStudy (lastdaywork):

I found the (generally accepted) formula to be inconsistent For xy>1 \[\tan^{-1}x+\tan^{-1}y=\tan^{-1}\frac{ x+y }{ 1-xy }+\pi\]

OpenStudy (lastdaywork):

@eliassaab

OpenStudy (ikram002p):

but still u need to know for tan^-1 a = - tan^-1 a so \(\large \tan^{-1}x+\tan^{-1}y=\tan^{-1}\frac{ x+y }{ 1-xy }+\pi\) \(\large -\tan^{-1}a-\tan^{-1}b=\tan^{-1}-\frac{(a+b) }{ 1-ab }+\pi\) \(\large -(\tan^{-1}a+\tan^{-1}b)=-\tan^{-1}\frac{(a+b) }{ 1-ab }+\pi\) \(\large -(\tan^{-1}a+\tan^{-1}b)=-(\tan^{-1}\frac{(a+b) }{ 1-ab }-\pi)\) does that help ?!

OpenStudy (lastdaywork):

So, what I posted as question is correct..right ?

OpenStudy (ikram002p):

yep since xy >1 x sighn = y sighn

OpenStudy (lastdaywork):

Thanxx :)

OpenStudy (ikram002p):

np :D

OpenStudy (anonymous):

Notice that \[ \tan \left(\tan ^{-1}(x)+\tan ^{-1}(y)\right)=\frac{\tan \left(\tan ^{-1}(x)\right)+\tan \left(\tan ^{-1}(x)\right)}{1-\tan \left(\tan ^{-1}(x)\right) \tan \left(\tan ^{-1}(x)\right)}=\frac{x+y}{1-x y} \\ \tan \left(\tan ^{-1}\left(\frac{x+y}{1-x y}\right)+\pi \right)=\tan \left(\tan ^{-1}\left(\frac{x+y}{1-x y}\right)\right)=\frac{x+y}{1-x y} \] and conclude

OpenStudy (anonymous):

Notice that \[ \tan(u)=\tan(v)\implies u=v+ k \pi \quad k=0,\pm1, \pm2 \pm 3 \]